OpenStudy (anonymous):

How do you solve this? x^4<=80

OpenStudy (kc_kennylau):

take the 4th root on both sides?

OpenStudy (anonymous):

But what do you do with the inequality signs?

OpenStudy (kc_kennylau):

ok i'd break it down to \(x^2\le\sqrt{80}\) and \(x^2\ge-\sqrt{80}\) which is rejected

OpenStudy (kc_kennylau):

Then I'd break \(x^2\le\sqrt{80}\) into \(x\le\sqrt[4]{80}\) and \(x\ge-\sqrt[4]{80}\).

OpenStudy (anonymous):

Thank you!

OpenStudy (kc_kennylau):

no problem :)